Hi all
sorry ,My english is not good , here question :
suppose we have sheet metal with 250mm*200mm*0.5mm and want to make "core like cylinder" with 100mm diameter & 200mm length, how many sheet metal do we need ?
thanks a lot.

Answer by mananth(16946)

Surface area of core = pi*D*L (neglect thickness)
=3.14*100*200
= 62,800 mm^2
..
sheet area = 250*200
=50,000 mm^2
..
sheet insufficient